# SPO: Sequential Monte Carlo Policy Optimisation

### v2

@@ 1 Introduction
The integration of reinforcement learning (RL) and neural-guided planning methods has recently achieved considerable success. Such methods effectively leverage planning during training via iterative imitation learning [5, 63, 16, 6]. They have demonstrated state of the art performance in applications ranging from chess [68] and matrix multiplication [22], to combinatorial optimisation and language modelling [40, 78]. However, one of the most commonly-used search-based policy improvement operator MCTS [14]: i) performs poorly on low budgets [29], ii) is inherently sequential limiting its scalability [47, 66], and iii) requires modifications to adapt to large or continuous action spaces [53, 37]. These limitations underscore the need for more scalable, efficient and generally applicable algorithms.
In this work, we introduce SPO: Sequential Monte Carlo Policy Optimisation, a model-based RL algorithm that utilises scalable sampled-based Sequential Monte Carlo planning for policy improvement. SPO is an approximate policy iteration algorithm, and formally ground it within the Expectation Maximisation (EM) framework.
SPO aims to leverage both breadth and depth of search in order to provide better estimates of the target distribution, derived via EM optimisation, compared to previous works. It combines the breadth used in MPO [1] and the depth used in V-MPO [69], while enforcing KL constraints on updates to ensure stable policy improvement. SPO attains strong performance across a range of benchmarks outperforming AlphaZero, a strong model-based expert iteration algorithm. We also benchmark against a leading Sequential Monte Carlo (SMC) algorithm from Piché et al. [60] and show that leveraging posterior estimates from SMC for policy improvement is central to strong performance.
Our work demonstrates that i) SPO outperforms baselines across both high-dimensional continuous control and challenging discrete planning tasks without modifications, ii) the sampling-based approach is inherently parallelisable, enabling the use of hardware accelerators to improve training speed. This provides a significant advantage over the sequential process of MCTS-based methods [14], which is particularly costly during training. iii) SPO empirically shows strong scaling behaviours, showing improved performance with additional search budget both during training and inference.
@@ 2 Related Work
There is an extensive history of prior works that frame control and RL as an inference problem [75, 42]. Expectation Maximisation [17, 24, 56] is a popular approach that seeks to directly optimise the evidence lower bound (ELBO) of the probability of optimality of a given policy. A number of model free methods implement EM, including RWR [58], REPS [59], MPO [1], V-MPO [69], and AWAC [55] each with different approaches to the E-step and M-step (summarised in Furuta et al. [25]). Model based algorithms such as MD-GPS [54] and AlphaZero [67] can also be viewed under the EM framework using LQR [3] and MCTS [7, 14] during the E-step respectively. Grill et al. [29] show MCTS is a form of regularised policy optimisation equivalent to optimising the ELBO. Parallelising MCTS with a virtual loss has been explored, however this often leads to performance degradation Liu et al. [47], increasing exploration and leading to out-of-distribution states that are difficult to evaluate Dalal et al. [15]. Model-based algorithms that iterate between policy improvement and projecting this improvement to the space of parameteriseable policies are often referred to as Expert iteration (ExIt) [67, 5] or Dual Policy Iteration [70]. Our work can be framed as ExIt, but differs by using Neural SMC for planning with an explicit KL regularisation on the improvement.
@@
Sequential Monte Carlo (SMC) [27, 39, 48], also referred to as particle filters [61, 4], is a particular inference method that can be applied to sample from an intractable distribution. Gu et al. [30] explore the parameterisation of the proposal in SMC using LSTMs [36], showing that posterior estimates from SMC can be used to train a parameterised proposal in non-control based tasks. Li et al. [45] demonstrate the same concept but using an MCMC sampler [26]. SMC can be applied to sample from distributions over trajectories in RL. For example, SMC-Learning [41] updates SMC weights using a Boltzmann exploration strategy, but unlike SPO it does not control policy improvement with KL regularisation or leverage neural approximators. Piché et al. [60] derive an SMC weight update to directly estimate the posterior over optimal trajectories. However this update requires the optimal value function, which they approximate using a SAC [33] value function. This approach requires high sample counts and does not leverage SMC posterior estimates for policy improvement, instead utilising a Gaussian distribution or SAC to train the proposal. CriticSMC [46] estimates the same posterior but utilises a soft action-value function to score particles [32], reducing the number of steps performed in the environment, enabling more efficient exploration and improved performance on lower SMC budgets compared with Piché et al. [60]. However their method is slower in wall clock time and uses a static proposal, therefore not performing iterative policy improvement.
@@ 3 Background
Sequential decision-making can be formalised under the Markov Decision Process (MDP) framework [62]. An MDP is a tuple $(\mathcal{S},\mathcal{A},\mathcal{T},r,\gamma,\mu)$ where, $\mathcal{S}$ is the set of possible states, $\mathcal{A}$ is the set of actions, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{P}(\mathcal{S})$ is the state transition probability function, $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function, $\gamma\in[0,1]$ is the discount factor, and $\mu$ is the initial state distribution. An agent interacts with the MDP using a policy a policy $\pi:\mathcal{S}\to\mathcal{P}(\mathcal{A})$, that associates to every state a distribution over actions. The quality of this interaction is quantified by the expected discounted return that the agent seeks to maximise $\pi^{*}\in\arg\max_{\pi\in\Pi}\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}\right],$ where $r_{t}=r(s_{t},a_{t})$ is the reward received at time $t$ and $\Pi$ is the set of all realisable policies. The value function $V^{\pi}(s_{t})=\mathbb{E}_{\pi}\left[\sum_{t=t}^{\infty}\gamma^{t}r_{t}\mid s_{t}\right]$ maps a state $s_{t}$ to the expected discounted sum of future rewards when behaving according to $\pi$. Similarly, the state-action value function $Q^{\pi}(s_{t},a_{t})=\mathbb{E}_{\pi}\left[\sum_{t=t}^{\infty}\gamma^{t}r_{t}\mid s_{t},a_{t}\right]$ maps a state $s_{t}$ to the expected discounted return, when taking the initial action $a_{t}$ and following $\pi$ thereafter.
@@
as $p(O=1|\tau)\propto\exp(\sum_{t}r_{t})$, $\alpha$ is a normalising constant, ensuring a valid probability distribution.
Expectation Maximisation (EM) [18] has been widely applied to solve the optimisation problem $\max_{\pi}\log p_{\pi}(\mathcal{O}=1)$ [17, 58, 59, 1, 69]. After deriving the lower bound $\mathcal{J}(q,\pi)$ on the objective in eq. 1 using an auxiliary non-parameteric distribution $q$, EM then performs coordinate ascent, iterating between optimising the bound with respect to $q$ (E-step) and with respect to the parameteric policy $\pi$ (M-step). This generates a sequence of policy pairs $\{(\pi_{0},q_{0}),(\pi_{1},q_{1}),\ldots,(\pi_{n},q_{n})\}$ such that in each step $i$ in the EM sequence, $\mathcal{J}(q_{i+1},\pi_{i+1})\geq\mathcal{J}(q_{i},\pi_{i})$. Viewed through a traditional policy improvement lens, the E-step corresponds to a policy evaluation phase where we perform rollouts, generating states with their associated estimates for $q$ and value estimates. The M-step corresponds to a policy improvement phase where $\pi$ and $V$ are updated.  [^1] Many related RL algorithms can be understood by the different approaches to either the E-step or M-step,see section F.1 for a summary of EM approaches.
@@ 4.1 Objective
We add an additional assumption to the lower bound objective defined in eq. 1 and assume that the auxiliary distribution $q$, defined over trajectories $\tau$, can be decomposed into individual state dependent distributions, i.e. $q(\tau)=\mu(s_{0})\prod_{t>0}q(a_{t}|s_{t})\mathcal{T}(s_{t+1}|s_{t},a_{t})$. We parameterise $\pi$ using $\theta$ which decomposes in the same way. This enables the objective to be written with respect to individual states instead of full trajectories. Multiplying by $\alpha$, the objective can then be written as follows:
@@
Previous works have explored this formulation where $p(\theta)$ is a prior over parameters of $\pi$ [34, 64, 1].
@@ 4.2 E-step
Equation 4 consists of balancing two objectives which we frame as a constrained maximisation problem, limiting the KL between $q$ and $\pi$ from exceeding a certain threshold:
@@
The following analytic solution to the constrained optimisation problem can then be derived using the Lagrangian multipliers method outlined in section G.1. Likewise $\eta^{*}$ is obtained by minimising the convex dual function eq. 7, and intuitively can be thought of as enforcing the KL constraint on $q$, preventing $q_{i}$ from moving to far from $\pi_{i}$, i.e.:
@@
where $V^{\bar{\pi}}(s)$ is an action independent baseline. $Q(s,a)-V(s)$ referred to as the advantage, where optimising for Q-values or advantages is equivalent [55], but has been demonstrated to outperform optimising Q-values directly [57, 69, 55]. This is also a natural update to perform as the policy improvement theorem [71] outlines that an update by policy iteration can be performed if at least one state-action pair has positive advantage and a non-zero probability of reaching such a state. Although we have a closed form solution for $q$, we do not have either the value function or action-value function in practice. In the next section we outline our approach to estimating this distribution.
Regularised Policy Optimisation: In our closed form solution to the optimisation eq. 6, we are required to solve for $\eta^{*}$ which ensures that the KL is constrained. This can be estimated by minimising the following objective [77], see section G.2 for derivation:
@@ 4.2.1 Estimating Target Distribution
Building upon previous work using Sequential Monte Carlo for trajectory distribution estimation [41, 60], we propose a novel approach that integrates SMC within the Expectation Maximisation framework to estimate the target distribution defined in Equation 6. Our SMC-based method enables sampling from $q_{i}$ for trajectories of length $h$, incorporating multiple predicted future states and rewards for a more accurate distribution estimation. A key property of estimating the target using SMC is that it uses both breadth (we initialise multiple particles to calculate importance weights) but also depth (leveraging future states and rewards) to estimate the distribution. In contrast, MPO evaluates several actions per state, focusing on breadth without depth, while V-MPO calculates advantages using n-step returns from trajectory sequences, leveraging depth but only for a single action sample. Our method combines both aspects, enhancing the accuracy of target distribution estimation, crucial for both action selection and effective policy improvement updates in the M-step. An outline of the SMC algorithm is provided in algorithm 1, with a visual representation in fig. 1.
SMC estimates the target by maintaining a set of particles $\{x^{(n)}\}_{n=1}^{N}$ each of which maintains an importance weight for a particular sample. Given calculated importance weights SMC, estimates the target distribution according to $\hat{q}_{i}{(\tau)}=\sum_{n=1}^{N}\bar{w}^{(n)}\delta_{x_{(n)}}(\tau)$ where $\bar{w}^{(n)}$ is the normalised importance sample weight for a particular sample. We next outline the sequential importance sampling weight update needed to estimate eq. 6. Since we can sample from $\pi(a_{t},s_{t},\theta_{i})$, we leverage this as our proposal distribution $\beta$. Given the target distribution $q$ can be decomposed into individual state dependent distributions, this leads to the following convenient weight update:
@@
Algorithm 1 outlines the method for estimating eq. 6 over $h$ planning steps for $N$ particles (line 3) using the advantage based weight update (line 7) in eq. 8. Once $h$ steps in the environment model have been performed in parallel, we marginalise all but the first actions as a sample based estimate of $q_{i}$ (line 14). The advantage function $A^{\bar{\pi}}$ is typically unknown, so we compute a 1-step estimate at each iteration. After $h$ steps, the importance weights leverage $h$-steps of observed states and rewards during planning and corresponding advantage estimates. Compared to tree-based methods such as MCTS, SMC does not require maintaining the full tree in memory. Instead, after each planning step, it only needs to retain the initial action, current state, and current particle weight.
Resampling Adaptive Search: To mitigate the issue of weight degeneracy [52], SPO conducts periodic resampling (lines 8-11). This involves generating a new set of particles from the existing set by duplicating some trajectories and removing others, based on the current particle weights. This process also enhances computational efficiency in estimating the target distribution. By resampling, we avoid updating weights for trajectories with low likelihood under the target, thereby reallocating computational resources to particles with high importance weights [20].

### v3

@@ 1 Introduction
The integration of reinforcement learning (RL) and neural-guided planning methods has recently achieved considerable success. Such methods effectively leverage planning during training via iterative imitation learning [7, 68, 18, 8]. Applying additional computation via planning generates improved policies which is then amortised [4] into a neural network policy. Using this policy as part of planning itself creates a powerful self-improvement cycle. They have demonstrated state of the art performance in applications ranging from chess [75] and matrix multiplication [24], to language modelling [61]. However, one of the most commonly-used search-based policy improvement operators, MCTS [16]: i) performs poorly on low budgets [33], ii) is inherently sequential limiting its scalability [51, 73], and iii) requires modifications to adapt to large or continuous action spaces [57, 41]. These limitations underscore the need for more scalable, efficient and generally applicable algorithms.
In this work, we introduce SPO: Sequential Monte Carlo Policy Optimisation, a model-based RL algorithm that utilises scalable sampled-based Sequential Monte Carlo planning for policy improvement. We formalise SPO as an approximate policy iteration algorithm, and show that it is formally grounded within the Expectation Maximisation (EM) framework.
SPO is unique by leveraging both breadth and depth of search in order to provide better estimates of the target distribution, derived via EM optimisation, compared to previous works. When viewed relative to popular algorithms it combines the breadth used in MPO [1] and the depth used in V-MPO [76], while enforcing KL constraints on updates to ensure stable policy improvement. SPO achieves strong performance across a range of benchmarks, outperforming AlphaZero, a leading model-based expert iteration algorithm. We also benchmark against a leading Sequential Monte Carlo (SMC) algorithm from Piché et al. [65] and show that leveraging posterior estimates from SMC for policy improvement is central to strong performance.
We demonstrate that i) SPO outperforms baselines across both high-dimensional continuous control and challenging discrete planning tasks without algorithmic modifications, and ii) the sampling-based approach is inherently parallelisable, enabling the use of hardware accelerators to improve training speed. This provides significant advantages over MCTS-based methods [16], whose sequential process can be particularly inefficient during training. 3) The explicit use of KL targeting leads to strong and stable policy improvement, reducing the need for a grid search over exploration hyperparameters in search. We find that SPO, empirically has strong scaling behaviours, with improved performance given additional search budget both during training and inference.
@@ 2 Related Work
There is an extensive history of prior works that frame control and RL as an inference problem [82, 45]. Expectation Maximization [19, 28, 60] is a popular approach to directly optimising the evidence lower bound (ELBO) of the probability of optimality of a given policy. A number of model free methods implement EM, including RWR [63], REPS [64], MPO [1], V-MPO [76], and AWAC [59] each with different approaches to the E-step and M-step (summarised in Furuta et al. [29]). Model based algorithms such as MD-GPS [58] and AlphaZero [74] can also be viewed under the EM framework using LQR [5] and MCTS [9, 16] during the E-step respectively. Grill et al. [33] show MCTS is a form of regularised policy optimisation equivalent to optimising the ELBO. Such model-based algorithms, that iterate between policy improvement using search, and projecting this improvement to the space of parameteriseable policies, are also often referred to as Expert iteration (ExIt) [74, 7] or Dual Policy Iteration [77]. Such approaches have also been applied to active inference [27] where MCTS has been used to estimate expected free energy, with this computation then amortised into a neural policy [25]. Our work can also be framed as ExIt, but differs by using Neural SMC for planning with an explicit KL regularisation on the improvement.
@@
Sequential Monte Carlo (SMC) [31, 43, 52], also referred to as particle filters [66, 6], is an inference method often employed to sample from intractable distributions. Gu et al. [34] explore the parameterisation of the proposal in SMC using LSTMs [40], showing that posterior estimates from SMC can be used to train a parameterised proposal in non-control based tasks. Li et al. [49] demonstrate the same concept but using an MCMC sampler [30]. SMC can be applied to sample from distributions over trajectories in RL. For example, SMC-Learning [44] updates SMC weights using a Boltzmann exploration strategy, but unlike SPO it does not control policy improvement with KL regularisation or leverage neural approximators. Piché et al. [65] derive an SMC weight update to directly estimate the posterior over optimal trajectories. However this update requires the optimal value function, which they approximate using a SAC [37] value function. This approach requires high sample counts and does not leverage SMC posterior estimates for policy improvement, instead utilising a Gaussian distribution or SAC to train the proposal. CriticSMC [50] estimates the same posterior but utilises a soft action-value function to score particles [36], reducing the number of steps performed in the environment, enabling more efficient exploration and improved performance on lower SMC budgets compared with Piché et al. [65]. However their method is slower in wall clock time and uses a static proposal, therefore not performing iterative policy improvement. A useful properly of SMC search methods is that it is inherently parallelisable. Parallelising MCTS with a virtual loss has been explored, however this often leads to performance degradation [51], increasing exploration and leading to out-of-distribution states that are difficult to evaluate [17].
@@ 3 Background
Sequential decision-making can be formalised under the Markov Decision Process (MDP) framework [67]. An MDP is a tuple $(\mathcal{S},\mathcal{A},\mathcal{T},r,\gamma,\mu)$ where, $\mathcal{S}$ is the set of possible states, $\mathcal{A}$ is the set of actions, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{P}(\mathcal{S})$ is the state transition probability function, $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function, $\gamma\in[0,1]$ is the discount factor, and $\mu$ is the initial state distribution. An agent interacts with the MDP using a policy $\pi:\mathcal{S}\to\mathcal{P}(\mathcal{A})$, that associates to every state a distribution over actions. We quantify the quality of a policy by the expected discounted return, that the agent seeks to maximise $\pi^{*}\in\arg\max_{\pi\in\Pi}\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}\right],$ where $r_{t}=r(s_{t},a_{t})$ is the reward received at time $t$ and $\Pi$ is the set of all realisable policies. The value function $V^{\pi}(s_{t})=\mathbb{E}_{\pi}\left[\sum_{t=t}^{\infty}\gamma^{t}r_{t}\mid s_{t}\right]$ maps a state $s_{t}$ to the expected discounted sum of future rewards when acting according to $\pi$. Similarly, the state-action value function $Q^{\pi}(s_{t},a_{t})=\mathbb{E}_{\pi}\left[\sum_{t=t}^{\infty}\gamma^{t}r_{t}\mid s_{t},a_{t}\right]$ maps a state $s_{t}$ to the expected discounted return, when taking the initial action $a_{t}$ and following $\pi$ thereafter.
@@
Since $p(O=1|\tau)\propto\exp(\sum_{t}r_{t})$ and $\alpha$ is a normalising constant, ensuring a valid probability distribution.
Expectation Maximisation (EM) [20] has been widely applied to solve the optimisation problem $\max_{\pi}\log p_{\pi}(\mathcal{O}=1)$ [19, 63, 64, 1, 76]. After deriving the lower bound $\mathcal{J}(q,\pi)$ on the objective in eq. 1 using an auxiliary non-parametric distribution $q$, EM then performs a coordinate ascent, iterating between optimizing the bound with respect to $q$ (E-step) and with respect to the parametric policy $\pi$ (M-step). This generates a sequence of policy pairs $\{(\pi_{0},q_{0}),(\pi_{1},q_{1}),\ldots,(\pi_{n},q_{n})\}$ such that in each step $i$ in the EM sequence, $\mathcal{J}(q_{i+1},\pi_{i+1})\geq\mathcal{J}(q_{i},\pi_{i})$. Viewed through a traditional policy improvement lens, the E-step corresponds to a policy evaluation phase where we perform rollouts, generating states with their associated estimates for $q$ and value estimates. The M-step corresponds to a policy improvement phase where $\pi$ and $V$ are updated. This step can also be thought of as amortising the probabilistic inference computation performed in the E-step, into a neural network forward pass operation.  [^1] Many related RL algorithms can be understood by the different approaches to either the E-step or M-step,see section F.1 for a summary of EM approaches.
@@ 4.1 Objective
We add an additional assumption to the lower bound objective defined in eq. 1 and assume that the auxiliary distribution $q$, defined over trajectories $\tau$, can be decomposed into individual state dependent distributions, i.e. $q(\tau)=\mu(s_{0})\prod_{t\geq 0}q(a_{t}|s_{t})\mathcal{T}(s_{t+1}|s_{t},a_{t})$. We parameterise $\pi$ using $\theta$ which decomposes in the same way. This enables the objective to be written with respect to individual states instead of full trajectories. Multiplying by $\alpha$, the objective can then be written as follows:
@@
Previous works have explored this formulation where $p(\theta)$ is a prior over the parameters of $\pi$ [38, 71, 1].
@@ 4.2 E-step
Equation 4 consists of balancing two objectives. To optimize this, we frame this as a constrained maximization problem, to avoid scaling problems, see section G.3 for further information. This objective limits the KL between $q$ and $\pi$ from exceeding a certain threshold:
@@
The following analytic solution to the constrained optimisation problem can then be derived using the Lagrangian multipliers method outlined in section G.1. Likewise $\eta^{*}$ is obtained by minimising the convex dual function eq. 7, and intuitively can be thought of as enforcing the KL constraint on $q$, preventing $q_{i}$ from moving to far from $\pi_{i}$ :
@@
where $V^{\bar{\pi}}(s)$ is an action independent baseline. $Q(s,a)-V(s)$ is referred to as the advantage, where optimising for Q-values vs advantages is equivalent [59]. Optimising with advantages however, has demonstrated to practically outperform optimising Q-values directly [62, 76, 59]. This is also a natural update to perform as the policy improvement theorem [78] outlines that an update by policy iteration can be performed if at least one state-action pair has positive advantage and a non-zero probability of reaching such a state. Although we have a closed form solution for $q$, we do not have either the value function or action-value function in practice. In the next section we outline our approach to estimating this distribution.
Regularised Policy Optimisation: In our closed form solution to the optimisation eq. 6, we are required to solve for $\eta^{*}$ which ensures that the KL is constrained. This can be calculated by minimising the following objective [84], see section G.2 for derivation:
@@
Practically, we estimate eq. 7 using a sample-based estimator according to the distribution of states from the replay buffer and stored values for $\pi(a|s,\theta_{i})$ and $A^{\bar{\pi}}(s,a)$.
@@ 4.2.1 Estimating Target Distribution
Building upon previous work using Sequential Monte Carlo for trajectory distribution estimation [44, 65], we propose a novel approach that integrates SMC within the Expectation Maximisation framework to estimate the target distribution defined in Equation 6. Our SMC-based method enables sampling from $q_{i}$ over trajectories of length $h$, incorporating multiple predicted future states and rewards for a more accurate distribution estimation. A key property of estimating the target using SMC is that it uses both breadth (we initialise multiple particles to calculate importance weights) but also depth (leveraging future states and rewards) to estimate the distribution. In contrast, MPO evaluates several actions per state, focusing on breadth without depth, while V-MPO calculates advantages using n-step returns from trajectory sequences, leveraging depth but only for a single action sample. Our method combines both aspects, enhancing the accuracy of target distribution estimation, crucial for both action selection and effective policy improvement updates in the M-step. An outline of the SMC algorithm is provided in algorithm 1, with a visual representation in fig. 1.
SMC estimates the target by maintaining a set of particles $\{x^{(n)}\}_{n=1}^{N}$ each of which maintains an importance weight for a particular sample. Given calculated importance weights SMC, estimates the target distribution according to $\hat{q}_{i}{(\tau)}=\sum_{n=1}^{N}\bar{w}^{(n)}\delta_{x_{(n)}}(\tau)$ where $\bar{w}^{(n)}$ is the normalised importance sample weight for a particular sample. We next outline the sequential importance sampling weight update needed to estimate eq. 6. Since we can sample from $\pi(a_{t},s_{t},\theta_{i})$, we leverage this as our proposal distribution $\beta$. Given the target distribution $q$ can be decomposed into individual state dependent distributions, we can define $p_{i}(\tau_{t}|\tau_{1:t-1})$ and $\beta(\tau_{t}|\tau_{1:t-1})$ as:
$p_{i}(\tau_{t}|\tau_{1:t-1})\propto\mathcal{T}(s_{t+1}|s_{t},a_{t})\pi_{i}(a_{t}|s_{t},\theta_{i})\exp\left(\frac{\bar{A_{i}}(a_{t},s_{t})}{\eta_{i}^{*}}\right)$ (8)
$\beta(\tau_{t}|\tau_{1:t-1})\propto\mathcal{T}_{\text{model}}(s_{t+1}|s_{t},a_{t})\pi_{i}(a_{t}|s_{t},\theta_{i})$ (9)
leading to the following convenient SMC weight update according to eq. 2:
@@
Algorithm 1 outlines the method for estimating eq. 6 over $h$ planning steps for $N$ particles (line 3) using the advantage based weight update (line 7) in eq. 10. Once $h$ steps in the environment model have been performed in parallel, we marginalise all but the first actions as a sample based estimate of $q_{i}$ (line 14).
@@
The advantage function $A^{\bar{\pi}}$ is typically unknown, so we compute a 1-step estimate at each iteration using the value function and observed environment rewards $\hat{A}(s_{t},a)=r_{t}+V^{\bar{\pi}}(s_{t+1})-V^{\bar{\pi}}(s_{t})$. Practically, we parameterise the value function $V$ using a neural network and train it using GAE [70] on true environment rollouts collected. After $h$ steps, the importance weights leverage $h$-steps of observed states and rewards during planning and corresponding advantage estimates. Compared to tree-based methods such as MCTS, SMC does not require maintaining the full tree in memory. Instead, after each planning step, it only needs to retain the initial action, current state, and current particle weight. It also does not require spending computation sending backward messages to update statistics of previous nodes everytime a new node is added to the tree.
Resampling Adaptive Search: To mitigate the issue of weight degeneracy [56], SPO conducts periodic resampling (lines 8-11). This involves generating a new set of particles from the existing set by duplicating some trajectories and removing others, based on the current particle weights. This process enhances computational efficiency in estimating the target distribution. By resampling, we avoid updating weights for trajectories with low likelihood under the target, thereby reallocating computational resources to particles with high importance weights [22].